# Quantum Multiple Eigenvalue Gaussian filtered Search: an efficient and versatile quantum phase estimation method

### 01 Feb 2024 v1

@@ Acknowledgment
This material is partially supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (Z.D.), by the Applied Mathematics Program of the US Department of Energy (DOE) Office of Advanced Scientific Computing Research under contract number DE-AC02-05CH1123 and the Simons Investigator program (L.L.), by the National Science Foundation under awards DMS-2011699 and DMS-2208163 (L.Y.), and by the Simons Institute for the Theory of Computing, through a Quantum Postdoctoral Fellowship (R.Z.). We thank the Institute for Pure and Applied Mathematics (IPAM) for its hospitality throughout the semester-long program “Mathematical and Computational Challenges in Quantum Computing” in Fall 2023 during which this work was initiated.
@@ Appendix A Proofs
where the inequality $\int_{x}^{\infty}e^{-\frac{1}{2}t^{2}}\mathrm{d}t\leq\sqrt{\frac{\pi}{2}}e^{-\frac{x^{2}}{2}}$ is used in the second last inequality. This proves (19). The proof of (20) and (21) is the same as the proof of [DDTL23, Appendix B.3 Lemma 4 eqn. (B8)], thus, we omit it here. ∎
@@
To prove the theorem, it suffices to show that when (26) and (27) hold, (14) holds. Fixed $1\leq k\leq|\mathcal{D}|$. Define $\mathcal{D}_{k_{-}}:=\{m\in\mathcal{D}:\lambda_{m}\in\mathcal{I}_{B,k}\}$, where $\mathcal{I}_{B,k}$ is the block interval defined in the algorithm. We note that $\mathcal{D}_{k_{-}}$ is the set of indices that have already been covered by the blocked set. Because $T=\Omega\left(\frac{\alpha}{\Delta_{\text{dom}}}\right)$, we find that each interval $\left[\boldsymbol{\theta}_{j}-\frac{\alpha}{T},\boldsymbol{\theta}_{j}+\frac{\alpha}{T}\right]$ contains exactly one dominant eigenvalue for all $11\leq j\leq k-$ 1 and ${\cal D}\backslash\mathcal{D}_{k_{-}}\neq\emptyset$. Then, to prove (14), it suffices to show that, if $\left|\boldsymbol{\theta}_{j}-\lambda_{m_{j}}\right|\leq\frac{Q}{T}$ for $j<k$iiiiiiiiiWhen $k=1$, we don’t need this condition., then there exists $m_{k}\in\mathcal{D}\setminus\mathcal{D}_{k_{-}}$ such that
@@ Appendix B Other quantum phase estimation algorithms
In this section, we give a brief summary of the previous quantum phase estimation algorithms that are tested in Section 4: MM-QCELS [DL23b], QPE (textbook version [NC00]), and ESPRIT [SHT22].
  • (MM-QCELS [DL23b]): The dataset used in MM-QCELS is similar to QMEGS (Section 2) and is also generated by Section 2. The main subroutine of MM-QCELS is called quantum complex exponential least squares (QCELS): Given a data set $\left\{(t_{n},Z_{n})\right\}^{N}_{n=1}$ generated from Algorithm 2, MM-QCELS obtains an estimate for the dominant eigenvalues by solving the following optimization problem:
@@
    Choosing $K=|\mathcal{D}|$ and a proper generated data set $\left\{(t_{n},Z_{n})\right\}^{N}_{n=1}$vvvIn [DL23b, Algorithm 2], the authors need to generate a sequence of data set $\left\{(t_{n},Z_{n})\right\}^{N}_{n=1}$ with different $T$ and $N$ to ensure Heisenberg limit scaling of the algorithm theoretically., [DL23b] shows that the solution $\{\theta^{*}_{k}\}^{K}_{k=1}$ of (48) is a good approximation to the set of dominant eigenvalues $\{\lambda_{m}\}_{m\in\mathcal{D}}$. MM-QCELS can reach Heisenberg limit scaling and small circuit depth when $p_{\mathrm{tail}}\ll 1$. However, the algorithm requires a spectral gap assumption, meaning $T=\Omega(1/\epsilon)>\Delta_{\text{dom}}^{-1}$, to ensure that the optimization problem can differentiate between the dominant eigenvalues.
  • (QPE (textbook version [NC00])): We provide a brief review of the textbook version of QPE. The quantum process of QPE involves a sequence of controlled time evolution operations $e^{-iH}$ on a state |$\ket{0^{d}}|\psi\rangle=\sum^{M-1}_{m=0}c_{m}\ket{0^{d}}\left|\psi_{m}\right\rangle$. Here, $\left|\psi_{m}\right\rangle$ represents the eigenstates associated with eigenvalues $\lambda_{m}$. The resulting quantum state from these operations, prior to applying the inverse Quantum Fourier Transform (QFT), is expressed as follows:
@@
    where $-N_{t}/2\leq k\leq N_{t}/2-1$, and $K_{N_{t}}$ is the squared and normalized Dirichlet kernel defined as $K_{N_{t}}(\theta)=\frac{\sin^{2}(\theta N_{t}/2)}{N_{t}^{2}\sin^{2}(\theta/2)}$. To simulate QPE classically, we sample this distribution $N_{QPE}$ times to obtain a set of samples $\{k_{i}\}^{N_{QPE}}_{i=1}$. The ground state energy can then be approximated as $\widetilde{\lambda}_{0}=\frac{2\pi\min{i}k_{i}}{N_{t}}$.
    It is noteworthy that, unlike other quantum phase estimation algorithms discussed in the numerical section, QPE is only capable of estimating the energy of the ground state, and it is not clear how to extend QPE to estimate multiple eigenvalues simultaneously.
  • (ESPRIT [SHT22]): The dataset used in MM-QCELS is similar to Section 2 and is also generated by Section 2. The algorithm of ESPRIT relies on the construction and manipulation of the Hankel matrix: Given $T>0$ and an odd integer $N>0$, we first set $t_{n}=n\tau$ for $0\leq n\leq N$, where $\tau=T/N$, and construct Hankel matrix $\mathrm{H}\in\mathbb{C}^{\frac{N+1}{2}\times\frac{N+1}{2}}$ with $\mathrm{H}_{i,j}=Z_{t_{i+j}}$. Here, $Z_{t_{n}}$ is generated by Section 2. We then find the singular value decomposition $\mathrm{H}=U\Sigma V^{\dagger}$ and define
@@
    Here, $U_{0}$ contains first $\frac{N-1}{2}$ rows and first $K$ columns of $U$ and $U_{1}$ contains the last $\frac{N-1}{2}$ rows and first $K$ columns of $U$. Finally, we find the eigenvalues $\{\mu_{k}\}^{K}_{k=1}$ of $U^{-1}_{1}U_{0}$ ($U^{-1}_{1}$ is the sudo-inverse of $U_{1}$) and define $\{\boldsymbol{\theta}_{k}=\mathrm{angle}(\mu_{k})/\tau\}^{K}_{k=1}$. According to the results of classical signal processing [LLF20], we can demonstrate that, when $K=|\mathcal{D}|$ and $T,N$ are chosen properly, $\{\boldsymbol{\theta}_{k}=\mathrm{angle}(\mu_{k})/\tau\}^{K}_{k=1}$ is a set that is close to the set of dominant eigenvalues $\{\lambda_{m}\}_{m\in\mathcal{D}}$.
    We would like to highlight that the original ESPRIT method falls short of achieving the Heisenberg limit scaling in the context of quantum phase estimation. In the original ESPRIT framework, the choice of $\tau=1$ is imperative to mitigate aliasing issuesviviviIt is not possible to differentiate between $\lambda_{k}$ and $\lambda_{k}+2\pi\tau$ in ESPRIT, as they produce the same data.. Consequently, this requires $N=T$, $T_{\max}=T$, and results in $T_{\mathrm{total}}=\Theta(T^{2}_{\max})$. However, according to the generalized uncertainty relation [BCM96], there exists a uniform complexity lower bound for phase estimation [GLM06], asserting that the square of the error is at least $\Omega\left(T_{\mathrm{total}}^{-1}T_{\max}^{-1}\right)$ in expectation. By combining $T_{\mathrm{total}}^{-1}T_{\max}^{-1}=\mathcal{O}(\epsilon^{2})$ and $T_{\mathrm{total}}=\Theta(T^{2}_{\max})$, we deduce $T_{\mathrm{total}}=\Omega(\epsilon^{-4/3})$. The crux of the matter, as deduced from the previous analysis, is that the necessary selection of $\tau=1$ to avoid aliasing poses a significant hurdle to ESPRIT in achieving the Heisenberg limit.
    More recently, [LNY23a] proposes a multilevel ESPRIT approach to circumvent aliasing issues without enforcing $\tau=1$. Notably,  [LNY23a] generates a sequence of datasets with carefully chosen values for $T$ and $N$, progressively refining the estimation of dominant eigenvalues. The successful application of multilevel techniques enables them to achieve Heisenberg-limited scaling and shorter circuit depth. In our numerical simulations presented in Section 4, for simplicity, we only consider the original ESPRIT and choose $N=T$ and $\tau=1$ to mitigate aliasing issues associated with ESPRIT.
@@ Appendix C Algorithm for integer-power setting
The monotonicity part is proved in [LNY23b, Lemma 2], where only some normalizing constants differ. Without loss of generality, we will prove the rest of the lemma assuming $x\geq 0$ since $\phi_{p}(x)$ is an even function. For the left part of (53), we need to notice that when $x=0$, the equality holds. Moreover, when $x\in[0,\pi]$, we have
@@
where we used the fact that $W<1$ and bounded the summation using integration. This proves (54). The proof of (55) is the same as the proof of [DDTL23, Appendix B.3 Lemma 4 eqn. (B8)], thus, we omit it here. ∎

### 26 Sep 2024 v2

@@ Acknowledgment
This material is partially supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (Z.D.), by the Applied Mathematics Program of the US Department of Energy (DOE) Office of Advanced Scientific Computing Research under contract number DE-AC02-05CH1123 and the Simons Investigator program (L.L.), by the National Science Foundation under awards DMS-2011699 and DMS-2208163 (L.Y.), and by DOE Grant No. DE-SC0024124 (R.Z.). We thank the Institute for Pure and Applied Mathematics (IPAM) for its hospitality throughout the semester-long program “Mathematical and Computational Challenges in Quantum Computing” in Fall 2023 during which this work was initiated.
@@ Appendix A Proofs
where the inequality $\int_{x}^{\infty}e^{-\frac{1}{2}t^{2}}\mathrm{d}t\leq\sqrt{\frac{\pi}{2}}e^{-\frac{x^{2}}{2}}$ is used in the second last inequality. This proves (19). The proof of (20) and (21) is the same as the proof of [22, Appendix B.3 Lemma 4 eqn. (B8)], thus, we omit it here. ∎
@@
To prove the theorem, it suffices to show that when (26) and (27) hold, (14) holds. Fixed $1\leq k\leq|\mathcal{D}|$. Define $\mathcal{D}_{k_{-}}:=\{m\in\mathcal{D}:\lambda_{m}\in\mathcal{I}_{B,k}\}$, where $\mathcal{I}_{B,k}$ is the block interval defined in the algorithm. We note that $\mathcal{D}_{k_{-}}$ is the set of indices that have already been covered by the blocked set. Because $T=\Omega\left(\frac{\alpha}{\Delta_{\text{dom}}}\right)$, we find that each interval $\left[\boldsymbol{\theta}_{j}-\frac{\alpha}{T},\boldsymbol{\theta}_{j}+\frac{\alpha}{T}\right]$ contains exactly one dominant eigenvalue for all $11\leq j\leq k-$ 1 and ${\cal D}\backslash\mathcal{D}_{k_{-}}\neq\emptyset$. Then, to prove (14), it suffices to show that, if $\left|\boldsymbol{\theta}_{j}-\lambda_{m_{j}}\right|\leq\frac{Q}{T}$ for $j<k$vvvWhen $k=1$, we don’t need this condition., then there exists $m_{k}\in\mathcal{D}\setminus\mathcal{D}_{k_{-}}$ such that
@@ Appendix B Other quantum phase estimation algorithms
In this section, we give a brief summary of the previous quantum phase estimation algorithms that are tested in Section 4: MM-QCELS [18], QPE (textbook version [1]), and ESPRIT [14].
  • (MM-QCELS [18]): The dataset used in MM-QCELS is similar to QMEGS (Section 2) and is also generated by Section 2. The main subroutine of MM-QCELS is called quantum complex exponential least squares (QCELS): Given a data set $\left\{(t_{n},Z_{n})\right\}^{N}_{n=1}$ generated from Algorithm 2, MM-QCELS obtains an estimate for the dominant eigenvalues by solving the following optimization problem:
@@
    Choosing $K=|\mathcal{D}|$ and a proper generated data set $\left\{(t_{n},Z_{n})\right\}^{N}_{n=1}$viiviiviiIn [18, Algorithm 2], the authors need to generate a sequence of data set $\left\{(t_{n},Z_{n})\right\}^{N}_{n=1}$ with different $T$ and $N$ to ensure Heisenberg limit scaling of the algorithm theoretically., [18] shows that the solution $\{\theta^{*}_{k}\}^{K}_{k=1}$ of (48) is a good approximation to the set of dominant eigenvalues $\{\lambda_{m}\}_{m\in\mathcal{D}}$. MM-QCELS can reach Heisenberg limit scaling and small circuit depth when $p_{\mathrm{tail}}\ll 1$. However, the algorithm requires a spectral gap assumption, meaning $T=\Omega(1/\epsilon)>\Delta_{\text{dom}}^{-1}$, to ensure that the optimization problem can differentiate between the dominant eigenvalues.
  • (QPE (textbook version [1])): We provide a brief review of QPE for ground state energy estimation. The quantum process involves a sequence of controlled time evolution operations $e^{-iH}$ on a state |$\ket{0^{d}}|\psi\rangle=\sum^{M-1}_{m=0}c_{m}\ket{0^{d}}\left|\psi_{m}\right\rangle$. Here, $\left|\psi_{m}\right\rangle$ represents the eigenstates associated with eigenvalues $\lambda_{m}$. The resulting quantum state from these operations, prior to applying the inverse Quantum Fourier Transform (QFT), is expressed as follows:
@@
    where $-N_{t}/2\leq k\leq N_{t}/2-1$, and $K_{N_{t}}$ is the squared and normalized Dirichlet kernel defined as $K_{N_{t}}(\theta)=\frac{\sin^{2}(\theta N_{t}/2)}{N_{t}^{2}\sin^{2}(\theta/2)}$. To simulate QPE classically, we sample this distribution $N_{QPE}$ times to obtain a set of samples $\{k_{i}\}^{N_{QPE}}_{i=1}$. The ground state energy can then be approximated as $\widetilde{\lambda}_{0}=\frac{2\pi\min_{i}k_{i}}{N_{t}}$.
    We note that the original textbook version of QPE algorithm in [1, Chapter 5.2] is designed to estimate general eigenvalues, not necessarily the ground state energy. The version presented above is a variation specifically designed for ground state energy estimation. There are also other variations of QPE, such as the Gaussian/Kaiser window-based QPE [37, 38], where the authors implement different resource states to produce a more concentrated kernel and reduce $T_{\rm total}$. In addition, it might be possible to extend QPE to estimate multiple eigenvalues simultaneously: For example, in the ideal case where $p_{\text{tail}}=0$ and $\{2\pi\lambda_{m}\}_{m\in\mathcal{D}}$ are finite-digit numbers. In this ideal scenario, when $N_{t}$ is sufficiently large, we have
    $P(k)=\sum_{m\in\mathcal{D}}\left|c_{m}\right|^{2}\delta_{0}\left(\frac{2\pi k}{N_{t}}-\lambda_{m}\right)\,.$
    Therefore, measuring the ancilla qubits $\mathcal{O}(\mathrm{poly}(K))$ times is sufficient to obtain all the eigenvalues exactly. In the non-ideal case, we need to design an efficient method to post-process the output of QPE and obtain accurate estimations of multiple eigenvalues. To the best of our knowledge, we are unaware of any such procedure that has clear complexity analysis. Thus, for simplicity, we mainly consider the simple version of QPE in our paper.
  • (ESPRIT [14]): The dataset used in ESPRIT is similar to Section 2 and is also generated by Section 2. The algorithm of ESPRIT relies on the construction and manipulation of the Hankel matrix: Given $T>0$ and an odd integer $N>0$, we first set $t_{n}=n\tau$ for $0\leq n\leq N$, where $\tau=T/N$, and construct Hankel matrix $\mathrm{H}\in\mathbb{C}^{\frac{N+1}{2}\times\frac{N+1}{2}}$ with $\mathrm{H}_{i,j}=Z_{t_{i+j}}$. Here, $Z_{t_{n}}$ is generated by Section 2. We then find the singular value decomposition $\mathrm{H}=U\Sigma V^{\dagger}$ and define
@@
    Here, $U_{0}$ contains first $\frac{N-1}{2}$ rows and first $K$ columns of $U$ and $U_{1}$ contains the last $\frac{N-1}{2}$ rows and first $K$ columns of $U$. Finally, we find the eigenvalues $\{\mu_{k}\}^{K}_{k=1}$ of $U^{-1}_{1}U_{0}$ ($U^{-1}_{1}$ is the pseudoinverse of $U_{1}$) and define $\{\boldsymbol{\theta}_{k}=\mathrm{angle}(\mu_{k})/\tau\}^{K}_{k=1}$. According to the results of classical signal processing [35], we can demonstrate that, when $K=|\mathcal{D}|$ and $T,N$ are chosen properly, $\{\boldsymbol{\theta}_{k}=\mathrm{angle}(\mu_{k})/\tau\}^{K}_{k=1}$ is a set that is close to the set of dominant eigenvalues $\{\lambda_{m}\}_{m\in\mathcal{D}}$.
    We would like to highlight that the original ESPRIT method falls short of achieving the Heisenberg limit scaling in the context of quantum phase estimation. In the original ESPRIT framework, the choice of $\tau=1$ is imperative to mitigate aliasing issuesviiiviiiviiiIt is not possible to differentiate between $\lambda_{k}$ and $\lambda_{k}+2\pi\tau$ in ESPRIT, as they produce the same data.. Consequently, this requires $N=T$, $T_{\max}=T$, and results in $T_{\mathrm{total}}=\Theta(T^{2}_{\max})$. However, according to the generalized uncertainty relation [39], there exists a uniform complexity lower bound for phase estimation [40], asserting that the square of the error is at least $\Omega\left(T_{\mathrm{total}}^{-1}T_{\max}^{-1}\right)$ in expectation. By combining $T_{\mathrm{total}}^{-1}T_{\max}^{-1}=\mathcal{O}(\epsilon^{2})$ and $T_{\mathrm{total}}=\Theta(T^{2}_{\max})$, we deduce $T_{\mathrm{total}}=\Omega(\epsilon^{-4/3})$. The crux of the matter, as deduced from the previous analysis, is that the necessary selection of $\tau=1$ to avoid aliasing poses a significant hurdle to ESPRIT in achieving the Heisenberg limit.
    More recently, [19] proposes a multilevel ESPRIT approach to circumvent aliasing issues without enforcing $\tau=1$. Notably,  [19] generates a sequence of datasets with carefully chosen values for $T$ and $N$, progressively refining the estimation of dominant eigenvalues. The successful application of multilevel techniques enables them to achieve Heisenberg-limited scaling and shorter circuit depth. In our numerical simulations presented in Section 4, for simplicity, we only consider the original ESPRIT and choose $N=T$ and $\tau=1$ to mitigate aliasing issues associated with ESPRIT.
@@ Appendix C Algorithm for integer-power setting
The monotonicity part is proved in [23, Lemma 2], where only some normalizing constants differ. Without loss of generality, we will prove the rest of the lemma assuming $x\geq 0$ since $\phi_{p}(x)$ is an even function. For the left part of (53), we need to notice that when $x=0$, the equality holds. Moreover, when $x\in[0,\pi]$, we have
@@
where we used the fact that $W<1$ and bounded the summation using integration. This proves (54). The proof of (55) is the same as the proof of [22, Appendix B.3 Lemma 4 eqn. (B8)], thus, we omit it here. ∎